The sum of two numbers is 23. Seven times the first number less five times the second number is equal to 5. What are the two numbers?
a + b = 23
7a + 5b = 5
S: not sure what to do after that

Answer by unlockmath(1688)

Hello,
With this:
a + b = 23
7a + 5b = 5
You almost had is set up correctly. The right way is:
a + b = 23
7a - 5b = 5
We can solve this two different ways; Substitution or elimination.Let's do substitution by setting the first equation as:
a=23-b and plug this into the second equation as:
7(23-b)-5b= 5
Now we have one variable (b) to solve:
161-7b-5b=5
Simplify and subtract 161 to get:
12b= -156
Divide by 12 and we get:
b=13
Now that we know what the value of b is we can plug this into the first or second equation to find the answer for "a". I'lllet you figure that one out.
Make sense?
RJ
www.math-unlock.com